# Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

### v1

@@ 2.2 Existing FJSSP Benchmarks
These design differences between the benchmark collections complicates the selection of reasonable test instances. For example, it is difficult to decide which instances of what collection should be chosen to evaluate an algorithm on a group of rather complex FJSSP instances. Especially on such instances, performance advantages of meta-heuristics over commercial MILP solvers could be expected. Of course, the answer could be to test the solvers on all available instances. Yet, it can often be useful to consider specific problem groups separately in order to focus on the individual strengths of a solver. In cases where the exact optimum is already known, i.e. for instances of relatively low complexity, the computing time required to achieve the optimum would be of greater interest. On the other hand, improvements of the solution quality are desirable as long as the exact optimum is unknown.
In this work, all $402$ distinct FJSSP instances of the ten test problem collections in Table 1 are considered. The table also indicates the number of instances (# instances) contained in each collections as well as the average number of operations per job (# avg. ops per job). The latter is within the range of $3$ and $19.5$, indicating the differences between the proposed benchmarking sources. Apart from the benchmark sources in Table 1, other collections with close relationships must be named. These are the problem collections in [3] and [4]. The design in [4] is limited to problem instances of the classical Job Shop Scheduling Problem (JSSP) and thus renounces the flexibility to run any of the jobs on any machine. Conversely, the instances in [3] allow parallel processing of some operations within a job. Taking these instances into account would therefore make it necessary to extend the models. Due to the special properties listed, both benchmark sets range outside the scope of this paper and are thus excluded. Conclusively, to the best of our knowledge, the benchmark sources referred to in Table 1 represent an extensive set of established test instances matching the definition of the FJSSP in Sec. 2.1.
@@
No. Source name # instances # avg. ops per job Reference
1 BehnkeGeiger 60 5,0 [5]
2 Brandimarte 15 8,6 [6]
3 ChambersBarnes 21 11,7 [7]
4 DPpaulli 18 19,5 [8]
5 Fattahi 20 3,0 [9]
6 Kacem 4 3,2 [10]
7 Hurink Edata 66 8,9 [11]
8 Hurink Sdata 66 8,9 [11]
9 Hurink Rdata 66 8,9 [11]
10 Hurink Vdata 66 8,9 [11]
@@
The first row of such a file commonly contains two (optionally three) numbers: the first value determines the number of jobs $n$ and the second one specifies the number of machines $m$ in the FJSSP instance. An optional 3rd value may indicate the average number of machines per operation. Although the latter is not necessary for the implementation of the problem instance. In the example above, the number of jobs is $n=10$, there are $m=10$ machine available, and each operations can on average be processed on $1$ machine.
@@
The remaining rows of the text file then represent the individual jobs $J_{i}$ $(i=1,...,n)$ of the problem instance. The first value in each row gives the number of operations $n_{i}$ to be executed in job $J_{i}$. The second number is the number of machines $\lvert\mathcal{M}_{i,1}\rvert$ available to process the first operation $O_{i,1}$. This is followed by a sequence of $\lvert\mathcal{M}_{i,j}\rvert$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$. Next, the data for the second operation $O_{i2}$ are provided. In this way, all operations of all jobs are specified accordingly.
In order to facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance are given in the Appendix. In addition to the key properties such as the total number of jobs ($n$), machines ($m$), or operations(N) of an instance, seven other instance-specific statistics are provided that allow for a rough intuition of the individual benchmark instance complexity. These include the minimum ($T_{\min}$), maximum ($T_{\max}$) and average processing times ($\bar{T}$) measured over all possible machine-operation combinations, as well as the corresponding standard deviation $\sigma_{T}$. Moreover, the average number of operations per job $\frac{\texttt{N}}{n}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided. The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to.  [12]
@@
It indicates the degree of resilience in the FJSSP setting [13]. In case, all operations of all jobs can be processed by any machine, $\beta=1$ and the FJSSP is called fully flexible. Otherwise, i.e. $10<\beta<$ 1, it is only partially-flexible and, consequently, harder to solve. The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{ijk}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}\lvert\mathcal{M}_{i,j}\rvert}\end{split}$ (3)
@@
Figure 3: Illustration of duration variety $dv$ and flexibility $\beta$ across all considered FJSSP instances.
@@
Figure 3 displays scatter plots of two characteristic features of all benchmark instances. This provides an first overview of the diversity of the individual instances and reveals where similarities coincide. Details on the individual characteristics of the considered benchmark sources can be found in the Appendix. This rough insight into the benchmark properties alone supports the demand for a targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers.
2.3 Recent Work on Benchmarking Worker Flexibility in FJSSP
@@
In order to retrieve benchmark instances including worker flexibility, the instances are usually specifically constructed for the given use case. In most cases, the constructed instances are based on subsets of commonly used benchmark instances [14][15][16]. In other cases, benchmark instances are randomly generated or constructed [16][17][18]. While the construction process is sometimes explained ([19][14]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems. The algorithm used to enable the conversion of all FJSSP benchmark instances was developed similarly to the algorithm used by [19].

### v2

@@ 2.2 Foundational FJSSP benchmark problem instances
These design differences between the benchmark collections complicates the selection of reasonable test instances. For example, it is difficult to decide which instances of what collection should be chosen to evaluate an algorithm on a group of rather complex FJSSP instances. Especially on such instances, performance advantages of meta-heuristics over commercial MILP solvers could be expected. Of course, the answer could be to test the solvers on all available instances. Yet, it can often be useful to consider specific problem groups separately in order to focus on the individual strengths of a solver. In cases where the exact optimum is already known, i.e. for instances of relatively low complexity, the computing time required to achieve the optimum would be of greater interest. On the other hand, improvements of the solution quality are desirable if the exact optimum is unknown.
In this work, all $402$ distinct FJSSP instances of the ten test problem collections in Table 1 are considered. The table also indicates the number of instances (# inst) contained in each collections as well as other average characteristics, e.g. the average number of jobs per instance $\bar{n}$, or the average number of machines $\bar{m}$. The average number of operations per job $\bar{\texttt{N}}/\bar{n}$ lies within the range of $2.95$ and $19.49$, indicating the differences between the proposed benchmarking sources. Apart from the benchmark sources in Table 1, other collections with close relationships must be named. These are the problem collections in [33] and [34]. The design in [34] is limited to problem instances of the classical Job Shop Scheduling Problem (JSSP) and thus renounces the flexibility to run any of the jobs on any machine. Conversely, the instances in [33] allow parallel processing of some operations within a job. Taking these instances into account would therefore make it necessary to extend the models. Due to the special properties listed, both benchmark sets range outside the scope of this paper and are thus excluded. Conclusively, to the best of our knowledge, the benchmark sources referred to in Table 1 represent an extensive set of established test instances matching the definition of the FJSSP in Sec. 2.1.
@@
No. Source name # inst $\bar{n}$ $\bar{\texttt{N}}$ $\nicefrac{{\bar{\texttt{N}}}}{{\bar{n}}}$ $\bar{m}$ $\bar{\beta}$ $\bar{dv}$ Ref.
1 BehnkeGeiger 60 45.00 225.00 5.0 40.0 0.316 0.009 [4]
2 Brandimarte 15 20.33 171.87 8.56 9.13 0.310 0.007 [35]
3 ChambersBarnes 21 13.33 158.33 11.67 13.667 0.089 0.007 [36]
4 DPpaulli 18 15.00 292.00 19.49 7.667 0.330 0.004 [37]
5 Fattahi 20 5.35 17.40 2.95 5.1 0.517 0.098 [38]
6 Kacem 4 9.75 31.75 3.16 8.0 1.0 0.042 [39]
7 Hurink EData 66 14.76 133.38 8.85 8.848 0.151 0.010 [40]
8 Hurink SData 66 14.76 133.38 8.85 8.848 0.131 0.010 [40]
9 Hurink RData 66 14.76 133.38 8.85 8.848 0.258 0.010 [40]
10 Hurink VData 66 14.76 133.38 8.85 8.848 0.476 0.010 [40]
Overall 402 18.89 150.11 8.54 13.510 0.281 0.014
@@
The first row of such a file commonly contains two (optionally three) numbers: the first value determines the number of jobs $n$ and the second one specifies the number of machines $m$ in the FJSSP instance. An optional third value may indicate the average number of machines per operation. Although the latter is not necessary for the implementation of the problem instance. In the example above, the number of jobs is $n=10$, there are $m=10$ machine available, and each operations can on average be processed on $1$ machine.
@@
The remaining rows of the text file then represent the individual jobs $J_{i}$ $(i=1,...,n)$ of the problem instance. The first value in each row gives the number of operations $n_{i}$ to be executed in job $J_{i}$. The second number is the number of machines $m_{i,j}$ available to process the first operation $O_{i,1}$. This is followed by a sequence of $m_{i,j}$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$. Next, the data for the second operation $O_{i,2}$ are provided. In this way, all operations of all jobs are specified accordingly.
To facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance from each source are given in a public repository on Github [41].
In addition to the key properties such as the total number of jobs ($n$), machines ($m$), or operations (N) of an instance, seven other instance-specific statistics are provided that allow for a rough intuition of the individual benchmark instance complexity. These include the minimum ($T_{\min}$), maximum ($T_{\max}$) and average processing times ($\bar{T}$) measured over all possible machine-operation combinations, as well as the corresponding standard deviation $\sigma_{T}$. Moreover, the average number of operations per job $\nicefrac{{\texttt{N}}}{{n}}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided. The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to [32].
@@
It indicates the degree of resilience in the FJSSP setting [42]. In case, all operations of all jobs can be processed by any machine, $\beta=1$ and the FJSSP is called fully flexible. Otherwise, i.e. $10<\beta<$ 1, it is only partially-flexible and, consequently, harder to solve. The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{i,j,k}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}m_{i,j}}\end{split}$ (3)
@@
Figure 3: An overview over the benchmark instances on a subset of problem characteristics to show the distribution and similarities between instances. The instances are marked with respect to their original source.
@@
Figure 3 displays scatter plots of two characteristic features of all benchmark instances. This provides a first overview of the diversity of the individual instances and reveals where similarities coincide. This illustration of benchmark properties supports the demand for targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers. Details on the particular characteristics of the considered benchmark sources can be found in the Github repository of the FJSSP(-W) benchmarking environment [41]. The description of the respective environment is provided in Sec. 4.
3 Worker Flexibility
Worker flexibility is a pivotal aspect of modern job scheduling, where the ability to adapt to varying worker skills and availability can significantly increase operational efficiency in production environments. This section addresses the Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W). It considers relevant benchmarking advances, provides a proper problem formulation, and discusses the design of the FJSSP-W instance set.
3.1 Recent work on FJSSP-W benchmarking
@@
To obtain benchmark instances that take into account the flexibility of the workforce, the instances are usually designed specifically for the respective application. In most cases, the constructed instances are based on subsets of frequently used benchmark instances [43, 23, 24]. Otherwise, benchmark instances are randomly generated or constructed [24, 44, 45]. While the construction process is sometimes explained ([46, 43]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems. The algorithm that enables the conversion of the FJSSP benchmark instances was developed similarly to the algorithm used by [47]. However, to make the process easier to reproduce, the method to derive the durations based on the workers has been adapted to sample from a range around the original duration and default parameters for all other parameters (e.g. amount of workers) are given.